OpenStudy (anonymous):

Is (1, 1) a solution to 3y = 2x + 1?

OpenStudy (anonymous):

yes it is its equal to zero

OpenStudy (anonymous):

try to substitute (1,1) into the equation...x=1 and y=1.... 3y=3(1)=3 2(1)+1=3 thus (1,1) is the solution...

OpenStudy (anonymous):

2x-3y+1 2(1)-3(1)+1 =0